OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 9 −x − y = −8 −8 + y −8 − y 8 + y 8 − y

OpenStudy (anonymous):

Well, what do you get when you isolate x in -x-y=-8

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

Well. step 1: add y to both sides step 2:divide both sides by -1

OpenStudy (anonymous):

-x -y= -8 1 1 -x = -7

OpenStudy (anonymous):

Wheres y? -x-y=-8 add y to both sides -x=-8-y divide both sides by -1 x=8+y

OpenStudy (anonymous):

x=8+y /1 /1

OpenStudy (anonymous):

We're finished, you substitute 8+y

OpenStudy (anonymous):

C